OpenStudy (anonymous):

How many pounds of an $83 per pound substance must be mixed with an $11 per pound substance to get 240 lbs of an alloy which will cost $28 per pound? Round to the nearest pound. 57 lbs 183 lbs 205 lbs 35 lbs

OpenStudy (jewlzme17):

x = lbs of $83 y = lbs of $11 You have two unknowns, so you need two equations x + y = 240 lbs & (83x + 11y)/240 = $28 or 83x + 11y = 6720 Change the first equation to y = 240 - x Now plug it into the second equation 83x + 11(240 - x) = 6720 so 83x + 2640 - 11x = 6720 so 72x = 4080 so x = 56.67, so the answer is a.